# Flow Matching with Gaussian Process Priors for Probabilistic Time Series Forecasting

### v1

@@ 4.1.2 Training Downstream Models
TSDiff $\underline{0.094}$±$\pm 0.007$ 0.010±$\pm 0.000$ $\mathbf{0.651}$±$\pm 0.036$ $0.042$±$\pm 0.014$ $0.634$±$\pm 0.009$ $0.235$±$\pm 0.006$ $0.392$±$\pm 0.011$ $0.369$±$\pm 0.027$
TSFlow Isotropic $\underline{0.094}$±$\pm 0.004$ $\underline{0.011}$±$\pm 0.000$ $\underline{0.691}$±$\pm 0.062$ $\mathbf{0.032}$±$\pm 0.003$ $\underline{0.608}$±$\pm 0.011$ $\underline{0.233}$±$\pm 0.001$ $\underline{0.345}$±$\pm 0.011$ $\underline{0.337}$±$\pm 0.005$
OU $0.098$±$\pm 0.008$ $\underline{0.011}$±$\pm 0.000$ $0.745$±$\pm 0.019$ $\underline{0.035}$±$\pm 0.002$ $0.618$±$\pm 0.009$ $0.235$±$\pm 0.003$ $0.350$±$\pm 0.005$ $0.342$±$\pm 0.011$
SE $0.096$±$\pm 0.003$ $\underline{0.011}$±$\pm 0.000$ $0.748$±$\pm 0.022$ $\underline{0.035}$±$\pm 0.003$ $0.616$±$\pm 0.003$ $0.238$±$\pm 0.002$ $0.351$±$\pm 0.007$ $0.348$±$\pm 0.010$
PE $\mathbf{0.090}$±$\pm 0.002$ $\underline{0.011}$±$\pm 0.000$ $0.812$±$\pm 0.043$ $\mathbf{0.032}$±$\pm 0.003$ $\mathbf{0.606}$±$\pm 0.008$ $\mathbf{0.229}$±$\pm 0.002$ $\mathbf{0.337}$±$\pm 0.006$ $\mathbf{0.332}$±$\pm 0.004$
@@ 4.2 Probabilistic Forecasting
SSSD $\underline{0.048}$±$\pm 0.001$ $\pm 0.001$ $\underline{0.274}$±$\pm 0.009$ $0.050$±$\pm 0.007$ $0.384$±$\pm 0.023$ $0.097$±$\pm 0.002$ $\underline{0.156}$±$\pm 0.007$ $\underline{0.209}$±$\pm 0.004$
@@
TSDiff $0.049$±$\pm 0.000$ $0.011$±$\pm 0.001$ $0.311$±$\pm 0.026$ $0.036$±$\pm 0.001$ $\underline{0.358}$±$\pm 0.020$ $0.098$±$\pm 0.002$ $0.172$±$\pm 0.005$ $0.221$±$\pm 0.001$
TSFlow-Cond. $\mathbf{0.045}$±$\pm 0.001$ $0.009$±$\pm 0.001$ $\mathbf{0.273}$±$\pm 0.004$ 0.025±$\pm 0.001$ 0.343±$\pm 0.002$ $\mathbf{0.083}$±$\pm 0.000$ $\mathbf{0.153}$±$\pm 0.001$ $0.227$±$\pm 0.000$
TSFlow-Uncond. $0.049$±$\pm 0.001$ $0.011$±$\pm 0.001$ $0.299$±$\pm 0.004$ $\underline{0.029}$±$\pm 0.001$ $0.464$±$\pm 0.004$ $\underline{0.089}$±$\pm 0.000$ $\mathbf{0.153}$±$\pm 0.002$ $0.279$±$\pm 0.007$
@@
Figure 4: Example forecasts and ground truth of TSFlow-Cond. of the first three time series in the test set of Traffic.
We test two variants of TSFlow: a conditional model with domain-specific Gaussian process regression prior (TSFlow-Cond., see Sec. 3.2), and an unconditional model with conditional prior sampling (TSFlow-Uncond., see Sec. 3.1.2). For simplicity, TSFlow-Cond. and TSFlow-Uncond. both use the Ornstein-Uhlenbeck kernel as it performed best in the forecasting setting (see Tab. 9). Tab. 3 shows that TSFlow is competitive with statistical and neural baselines and achieves state-of-the-art results on various datasets. On 6/8 datasets, TSFlow attains the best scores with up to 14% improvement to the second-best method. Furthermore, TSFlow-Cond. outperforms the diffusion-based baselines CSDI, SSSD, (Biloš et al., 2023), and TSDiff on 7/8 datasets while requiring fewer NFEs. Finally, we observe that, while TSFlow-Uncond. is competitive with other baselines, it is slightly inferior to the conditional version of TSFlow, showing that the conditional training is an effective method to specialize TSFlow for forecasting. We show example forecasts of TSFlow-Cond. on the Traffic dataset in Fig. 4 and for Electricity and Solar in Sec. A.7. Furthermore, we provide an ablation study of TSFlow’s different components in Sec. B.1.
@@ 5 Related Work
Diffusion models have been applied to several fields and achieved state-of-the-art performance (Ho et al., 2020; Hoogeboom et al., 2022; Lienen et al., 2024). Various works have demonstrated the effectiveness of unconditional models in conditional tasks through diffusion guidance (Epstein et al., 2023; Dhariwal & Nichol, 2021; Bansal et al., 2023; Nichol et al., 2021; Avrahami et al., 2022), solving inverse problems (Kawar et al., 2022), or iterative resampling (Lugmayr et al., 2022). Conditional Flow Matching (Lipman et al., 2022) is a recent approach proposed as an alternative to and generalization of diffusion models. This framework has been extended to incorporate couplings between data and prior samples, and trained to solve the dynamic optimal transport problem (Tong et al., 2023), and has been further adapted to general geometries (Chen & Lipman, 2023). Albergo et al. (2023) demonstrated the applicability of data-dependent couplings for stochastic interpolants to improve in-painting and super-resolution image generation. Lastly, recent work shows the potential for conditional generation by solving an optimization problem and differentiating through the flow (Ben-Hamu et al., 2024).
@@
Various generative models have been adapted to time series modeling, including Generative Adversarial Networks (GANs) (Yoon et al., 2019), normalizing flows (Rasul et al., 2020; Alaa et al., 2020), and Variational Autoencoders (VAEs) (Desai et al., 2021). Recent works have successfully applied diffusion models to time series. The first approach, TimeGrad (Rasul et al., 2021), applies a diffusion model on top of an LSTM to perform autoregressive multivariate time series forecasting, which was later extended by Biloš et al. (2023) to settings with continuous time and non-isotropic noise distributions. Note that unlike TSFlow, Biloš et al. (2023) neither use optimal transport paths nor couplings. CSDI (Tashiro et al., 2021) and SSSD (Alcaraz & Strodthoff, 2022) perform probabilistic forecasting and imputation via conditional diffusion models. While CSDI uses transformer layers for their backbone, SSSD makes use of S4 layers (Gu et al., 2021) to model temporal dynamics. Lastly, TimeDiff (Shen & Kwok, 2023) explores conditioning mechanisms, and TSDiff (Kollovieh et al., 2023) introduced an unconditional diffusion model that allows conditioning during inference time through diffusion guidance.
@@ 6 Conclusion
While TSFlow has demonstrated remarkable performance across multiple benchmarks and reduced the computational costs compared to its diffusion-based predecessors, it has been tested only on univariate time series. Future work could extend this approach to multivariate time series by leveraging multivariate Gaussian processes that capture dependencies across dimensions. Additionally, exploring source distributions with intractable likelihoods, such as neural forecasting methods, might further improve the performance.
@@
To ensure the reproducibility of our results, we provide a detailed description of our experimental setup, including the benchmark datasets and evaluation metrics, in Sec. 4. Additionally, we outline the hyperparameters used in Sec. A.2 and include pseudocode for our methods in Sec. A.8.
@@ A.2 Hyperparameters and training details
We trained both the conditional and unconditional variants of TSFlow using the Adam optimizer with a learning rate of $1\times 10^{-3}$, applying gradient clipping with a threshold of 0.5. The conditional model was trained for 400 epochs, while the unconditional model was trained for 1000 epochs. Each epoch consists of 128 batches, with each batch containing 64 sequences. To stabilize the training process, we employed an Exponential Moving Average (EMA) of the model parameters with a decay rate of $\beta=0.9999$. To parametrize the Gaussian processes, we choose $\ell=\sqrt{\nicefrac{{1}}{{2}}}$, $\ell=1$, and $\ell=\sqrt{2}$, for the SE, OU, and PE kernels, respectively, simplifying these even further.
@@
The architecture of TSFlow consists of three residual blocks with a hidden dimension of 64, resulting in approximately 176,000 trainable parameters. We encode timesteps using 64-dimensional sinusoidal embeddings (Vaswani et al., 2017).
To solve the ordinary differential equation (ODE), we utilized an Euler solver with 32 steps. The unconditional version, TSFlow-Uncond., uses 16 additional NFEs due to the additional conditional prior sampling, which involves 4 iterations with 4 steps each. The unconditional version in Tab. 2 uses 4 steps. TSFlow-Uncond. performs four iterations to sample from the prior distribution with a step size of $\tau=0.001$ and a noise scale of $0.01$. The guidance scale parameter is selected between $4$ and $6$ based on the validation set’s performance. Finally, we choose $\kappa$ in the guidance score uniformly between $0.1$ and $0.9$. An overview of the hyperparameters and training details is provided in Table 5.
@@
$\mathbf{x}_{\mathbf{p}}^{\mathrm{norm}}=\frac{\mathbf{x}_{\mathbf{p}}}{\mathrm{mean}(\mathbf{x}_{\mathrm{hist}})},$
@@
A.3 Baselines
We utilize the official implementations provided by the respective authors for CSDI, TSDiff, and Biloš et al. (2023). Although the original CSDI paper focuses on multivariate time series experiments, their codebase supports univariate forecasting, allowing us to use it without any modifications with their recommended hyperparameters and training setup. For SSSD, we made slight adjustments because their experiments primarily targeted multivariate datasets, whereas our focus is on univariate settings. Specifically, to avoid overfitting, we set the hidden dimension to 64 and the number of residual layers to 3, aligning the network sizes with those of TSDiff and TSFlow. For PatchTST, ARIMA, and ETS, we use AutoGluon (Shchur et al., 2023). For the remaining baselines, we employ the publicly available implementations in GluonTS (Alexandrov et al., 2020) with their recommended hyperparameters.
A.4 Unconditional Priors
Since the observations occur at discrete time steps, i.e., $t_{i}=i$ for $i=0,\dots,L-1$, we normalize them to align with the periodicity of the dataset. More specifically, we adjust the time steps using the formula:
$t_{i}^{\mathrm{norm}}=\frac{t_{i}\cdot\pi}{p},$ (15)
where $p$ represents the frequency length of the time series; for instance, 24 for hourly data and 30 for daily data. This normalization scales the time steps so that one full period corresponds to $\pi$, ensuring that the periodic kernel accurately captures the cyclical patterns in the data. Examples of the different processes for various kernels are illustrated in Fig. 5. To avoid numerical issues such as ill-conditioned covariance matrices, we extend each kernel $K$ with a white noise kernel $K_{\mathrm{white}}(t,t^{\prime})=\delta(t-t^{\prime})$ to $K^{\prime}(t,t^{\prime})=K(t,t^{\prime})+K_{\mathrm{white}}(t,t^{\prime})$. Adding this white noise ensures that the covariance matrix is positive definite and improves numerical stability during computations.

### v2

@@ 4.1.2 Training Downstream Models
TSDiff $0.094$±$\pm 0.007$ $\mathbf{0.010}$±$\pm 0.000$ $\mathbf{0.651}$±$\pm 0.036$ $0.042$±$\pm 0.014$ $0.634$±$\pm 0.009$ $0.235$±$\pm 0.006$ $0.392$±$\pm 0.011$ $0.369$±$\pm 0.027$
TSFlow Isotropic $\underline{0.089}$±$\pm 0.002$ $\mathbf{0.010}$±$\pm 0.000$ $\underline{0.684}$±$\pm 0.047$ $\underline{0.031}$±$\pm 0.001$ $\underline{0.607}$±$\pm 0.003$ $\underline{0.232}$±$\pm 0.003$ $0.360$±$\pm 0.003$ $\underline{0.349}$±$\pm 0.005$
OU $0.096$±$\pm 0.007$ $\underline{0.011}$±$\pm 0.000$ $0.722$±$\pm 0.046$ $0.032$±$\pm 0.001$ $0.616$±$\pm 0.002$ $0.237$±$\pm 0.001$ $0.352$±$\pm 0.007$ $0.361$±$\pm 0.003$
SE $0.102$±$\pm 0.007$ $\underline{0.011}$±$\pm 0.000$ $0.726$±$\pm 0.056$ $0.032$±$\pm 0.001$ $0.621$±$\pm 0.001$ $0.241$±$\pm 0.003$ $\underline{0.351}$±$\pm 0.008$ $0.365$±$\pm 0.003$
PE $\mathbf{0.083}$±$\pm 0.006$ $0.012$±$\pm 0.000$ $0.900$±$\pm 0.082$ $\mathbf{0.029}$±$\pm 0.001$ $\mathbf{0.590}$±$\pm 0.003$ $\mathbf{0.225}$±$\pm 0.003$ $\mathbf{0.320}$±$\pm 0.006$ $\mathbf{0.339}$±$\pm 0.005$
@@ 4.2 Probabilistic Forecasting
SSSD $\underline{0.048}$±$\pm 0.001$ $\pm 0.001$ $\mathbf{0.274}$±$\pm 0.009$ $0.050$±$\pm 0.007$ $0.384$±$\pm 0.023$ $0.097$±$\pm 0.002$ $0.156$±$\pm 0.007$ $0.209$±$\pm 0.004$
@@
TSDiff $0.049$±$\pm 0.000$ $0.011$±$\pm 0.001$ $0.311$±$\pm 0.026$ $0.036$±$\pm 0.001$ $0.358$±$\pm 0.020$ $0.098$±$\pm 0.002$ $0.172$±$\pm 0.005$ $0.221$±$\pm 0.001$
TSFlow-Cond. (ISO) $\mathbf{0.045}$±$\pm 0.001$ $\underline{0.008}$±$\pm 0.000$ $0.303$±$\pm 0.004$ $0.034$±$\pm 0.001$ $0.350$±$\pm 0.016$ $\mathbf{0.082}$±$\pm 0.000$ $\mathbf{0.154}$±$\pm 0.002$ $\underline{0.208}$±$\pm 0.000$
TSFlow-Cond. (OU) $\mathbf{0.045}$±$\pm 0.000$ $\underline{0.008}$±$\pm 0.000$ $0.288$±$\pm 0.004$ $\underline{0.028}$±$\pm 0.008$ $0.344$±$\pm 0.006$ $\mathbf{0.082}$±$\pm 0.000$ $\mathbf{0.154}$±$\pm 0.002$ $\mathbf{0.207}$±$\pm 0.001$
TSFlow-Cond. (SE) $\mathbf{0.045}$±$\pm 0.001$ $\underline{0.008}$±$\pm 0.000$ $0.296$±$\pm 0.002$ $\mathbf{0.027}$±$\pm 0.006$ $\underline{0.343}$±$\pm 0.011$ $\underline{0.083}$±$\pm 0.000$ $\underline{0.155}$±$\pm 0.002$ $0.211$±$\pm 0.003$
TSFlow-Cond. (PE) $\mathbf{0.045}$±$\pm 0.001$ $\underline{0.008}$±$\pm 0.000$ $\underline{0.278}$±$\pm 0.002$ $0.033$±$\pm 0.005$ $\mathbf{0.339}$±$\pm 0.005$ $\mathbf{0.082}$±$\pm 0.000$ $\underline{0.155}$±$\pm 0.003$ $0.211$±$\pm 0.001$
TSFlow-Uncond. $0.049$±$\pm 0.001$ $0.011$±$\pm 0.001$ $0.287$±$\pm 0.005$ $0.032$±$\pm 0.001$ $0.396$±$\pm 0.005$ $0.086$±$\pm 0.000$ $\mathbf{0.154}$±$\pm 0.002$ $0.291$±$\pm 0.005$
@@
Figure 2: Example forecasts and ground truth of TSFlow-Cond. (OU) of the first three time series in the test set of Traffic.
We evaluate two variants of TSFlow: a conditional model that uses domain-specific Gaussian process regression prior (TSFlow-Cond., see Sec. 3.2), and an unconditional model that applies conditional prior sampling and guidance (TSFlow-Uncond., see Sec. 3.1.2 and 3.1.3). Tab. 3 shows that TSFlow is competitive with statistical and neural baselines and achieves state-of-the-art results on various datasets. On 6/8 datasets, TSFlow attains the best scores with up to 14% improvement to the second-best method. Furthermore, TSFlow-Cond. outperforms or matches the diffusion-based baselines CSDI, SSSD, (Biloš et al., 2023), and TSDiff on 7/8 datasets while requiring fewer NFEs. Finally, we observe that, while TSFlow-Uncond. is competitive with other baselines, it is slightly inferior to the conditional versions of TSFlow, showing that the conditional training is an effective method to specialize TSFlow for forecasting. Furthermore, while all kernels yield similar results, the Ornstein-Uhlenbeck prior slightly improves over the others. We show example forecasts of TSFlow-Cond. on the Traffic dataset in Fig. 2 and for Electricity and Solar in Sec. A.9. Furthermore, we provide an ablation study of TSFlow’s different components in Sec. B.2 and extend the forecasting setting in Sec. B.7, B.6 and B.8.
@@ 5 Related Work
Diffusion models have been applied to several fields and achieved state-of-the-art performance (Ho et al., 2020; Hoogeboom et al., 2022; Lienen et al., 2024). Various works have demonstrated the effectiveness of unconditional models in conditional tasks through diffusion guidance (Epstein et al., 2023; Dhariwal & Nichol, 2021; Bansal et al., 2023; Nichol et al., 2021; Avrahami et al., 2022), solving inverse problems (Kawar et al., 2022), or iterative resampling (Lugmayr et al., 2022). Conditional Flow Matching (Lipman et al., 2022) is a recent approach proposed as an alternative to and generalization of diffusion models. This framework has been extended to incorporate couplings between data and prior samples, and trained to solve the dynamic optimal transport problem (Tong et al., 2023). Albergo et al. (2023) demonstrated the applicability of data-dependent couplings for stochastic interpolants to improve in-painting and super-resolution image generation. Lastly, recent work shows the potential for conditional generation by solving an optimization problem and differentiating through the flow (Ben-Hamu et al., 2024).
@@
Various generative models have been adapted to time series modeling, including Generative Adversarial Networks (GANs) (Yoon et al., 2019), normalizing flows (Rasul et al., 2020; Alaa et al., 2020), and Variational Autoencoders (VAEs) (Desai et al., 2021). Recent works have successfully applied diffusion models to time series. The first approach, TimeGrad (Rasul et al., 2021), applies a diffusion model on top of an LSTM to perform autoregressive multivariate time series forecasting, which was later extended by Biloš et al. (2023) to settings with continuous time and non-isotropic noise distributions. Note that unlike TSFlow, Biloš et al. (2023) neither use optimal transport paths nor couplings. CSDI (Tashiro et al., 2021) and SSSD (Alcaraz & Strodthoff, 2022) perform forecasting and imputation via conditional diffusion models. While CSDI uses transformer layers, SSSD makes use of S4 layers (Gu et al., 2021) to model temporal dynamics. Lastly, TimeDiff (Shen & Kwok, 2023) explores conditioning mechanisms, and TSDiff (Kollovieh et al., 2023) introduced an unconditional diffusion model conditioned through diffusion guidance similar to our unconditional model.
Additionally, Tamir et al. (2024) model dynamical systems using conditional flow matching (CFM), where the time series and ODE share the same time dimension. In contrast, TSFlow operates on two orthogonal time dimensions, one for the generative process and another within the time series. Moreover, Kerrigan et al. (2023) extend flow-matching to infinite-dimensional spaces using Gaussian processes, but their focus differs from ours. While they construct conditional Gaussian measure paths, TSFlow directly parameterizes the joint data and prior distribution.
@@ 6 Conclusion
While TSFlow achieves state-of-the-art performance across multiple benchmarks and reduces the computational costs compared to its diffusion-based predecessors, it has primarily been tested on univariate time series. Furthermore, different datasets and tasks, e.g., unconditional and conditional generation, require kernel selection, which should be optimized on the validation set. Future work could extend our approach to multivariate time series by incorporating multivariate Gaussian processes that capture dependencies across dimensions. Additionally, TSFlow allows for arbitrary source distributions, enabling the use of more involved priors, such as those based on data statistics (Park et al., 2024) or neural forecasting methods with intractable likelihoods. Exploring such distributions could further enhance performance.
@@
To ensure the reproducibility of our results, we provide a detailed description of our experimental setup, including the benchmark datasets and evaluation metrics, in Sec. 4 and A.1. Additionally, we outline the hyperparameters used in Sec. A.2 and include the pseudocodes for our unconditional and conditional model in Sec. A.10.
@@ A.2 Hyperparameters and training details
[…]
@@
The architecture of TSFlow is shown in Fig. 3 and consists of three residual blocks with a hidden dimension of 64, resulting in approximately 176,000 trainable parameters. We encode timesteps using 64-dimensional sinusoidal embeddings (Vaswani et al., 2017).
To solve the ordinary differential equation (ODE), we utilized an Euler solver with 32 steps. The unconditional version, TSFlow-Uncond., uses 16 additional NFEs due to the additional conditional prior sampling, which involves 4 iterations with 4 steps each. The unconditional version in Tab. 2 uses 4 steps. TSFlow-Uncond. performs four iterations to sample from the prior distribution with a step size of $\eta=0.005$ and a noise scale of $0.5$. The guidance scale parameter is selected between 8 and 32 based on the validation set’s performance. Finally, we choose $\kappa$ in the guidance score uniformly between $0.1$ and $0.9$. An overview of the hyperparameters and training details is provided in Table 5.
@@
$\mathbf{x}_{p}^{\mathrm{norm}}=\frac{\mathbf{x}_{p}}{\mathrm{mean}(\mathbf{x}_{\mathrm{hist}})},$
@@
A.3 Asymmetric Laplace Distribution
In Sec. 3.1.2 and 3.1.3, we parameterize the guidance score using an Asymmetric Laplace Distribution (ALD). The probability density function (PDF) is defined with a location parameter $m$, a scale parameter $\lambda$, and a quantile $\kappa$:
$p(x;m,\lambda,\kappa)\propto\exp\bigg{(}-\frac{1}{\lambda}\max\{\kappa\cdot(x-m),(\kappa-1)\cdot(x-m)\}\bigg{)}.$ (15)
By setting $\lambda=1$, the guidance score from Eq. 12 simplifies to the quantile loss:
$\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}^{p}\mid\mathbf{x}_{t})=\max\{\kappa\cdot(\mathbf{y}^{p}-\phi_{{\bm{\theta}},1}(\mathbf{x}_{t})),(\kappa-1)\cdot(\mathbf{y}^{p}-\phi_{{\bm{\theta}},1}(\mathbf{x}_{t}))\},$ (16)
which corresponds to a weighted $\ell_{1}$ loss. In practice, we select $\kappa$ uniformly from the range $(0.1,0.9)$ to ensure coverage across different quantiles.
A.4 Baselines
We utilize the official implementations provided by the respective authors for CSDI, TSDiff, and Biloš et al. (2023). Although the original CSDI paper focuses on multivariate time series experiments, their codebase supports univariate forecasting, allowing us to use it without any modifications with their recommended hyperparameters and training setup. As SSSD targeted primarily multivariate datasets, we adjusted their model to the univariate case. Specifically, to avoid overfitting, we set the hidden dimension to 64 and the number of residual layers to 3, aligning the network sizes with those of TSDiff and TSFlow. For PatchTST, ARIMA, and ETS, we use AutoGluon (Shchur et al., 2023). For the remaining baselines, we employ the publicly available implementations in GluonTS (Alexandrov et al., 2020) with their recommended hyperparameters.
A.5 Metrics
Linear Predictive Score (LPS)
To assess the performance of the unconditional models in downstream forecasting, we use the Linear Predictive Score (LPS) as proposed by Kollovieh et al. (2023). The LPS is computed as the test CRPS of a linear ridge regression model trained on synthetic samples generated by the model. Specifically, the linear model is trained to map the observed past of a time series, $\mathbf{y}^{p}\in\mathbb{R}^{L^{p}}$, to its future values $\mathbf{y}^{f}\in\mathbb{R}^{L^{f}}$. The linear model is regressed against the true future values, using the past $\mathbf{y}^{p}$ as input, allowing us to measure how well the synthetic samples capture the relationship between past and future. Since the linear ridge regression can be fit in closed form, it remains robust to random initialization. Following Kollovieh et al. (2023), we generate 10,000 synthetic samples to fit the linear model.
A.6 Unconditional Priors
Since the observations occur at discrete time steps, i.e., $\tau_{i}=i$ for $i=0,\dots,L-1$, we normalize them to align with the periodicity of the dataset. More specifically, we adjust the time steps using the formula:
$\tau_{i}^{\mathrm{norm}}=\frac{\tau_{i}\cdot\pi}{p},$ (17)
where $p$ represents the frequency length of the time series; for instance, 24 for hourly data and 30 for daily data. This normalization scales the time steps so that one full period corresponds to $\pi$, ensuring that the periodic kernel accurately captures the cyclical patterns in the data. Examples of the different processes for various kernels are illustrated in Fig. 4. To avoid numerical issues such as ill-conditioned covariance matrices, we extend each kernel $K$ with a white noise kernel $K_{\mathrm{white}}(\tau,\tau^{\prime})=\delta(\tau-\tau^{\prime})$ to $K^{\prime}(\tau,\tau^{\prime})=K(\tau,\tau^{\prime})+K_{\mathrm{white}}(\tau,\tau^{\prime})$. Adding this white noise ensures that the covariance matrix is positive definite and improves numerical stability during computations.